# Multimodal Referring Segmentation: A Survey

### v1

@@ 1 Introduction
Multimodal referring segmentation [1, 2, 3, 4, 5, 6, 7] aims to segment the target object in a visual scene of image [2, 3], video [1, 8], or 3D [9, 7] according to a referring expression, such as free-form text or audio. For example, as shown in Fig. 1(b), given the text referring expression “The bird flying away”, the model is expected to segment and track the described target object in the video. This task presents a fundamental and challenging problem in multimodal understanding and supports a wide range of practical applications, such as image/video editing [10, 11], robotics [12], autonomous driving [13], etc. Because of its significant potential in practical applications, multimodal referring segmentation has received growing attention in recent years, as shown in Fig. 3.
@@ 2.1.1 A Unified Formulation
where $f$ is typically instantiated as a neural network. The input space $\mathcal{X}$ consists of two components: the visual input $\mathcal{V}$ (e.g., image, silent video, auditory video, or 3D data), and the referring signal ${E}$ (e.g., text, audio, etc.) that specifies the target object(s) of interest. The output space $\mathcal{Y}$ consists of segmentation mask(s) of the referred entities within $\mathcal{V}$. Based on this unified formulation, we construct a comprehensive taxonomy of referring segmentation tasks and formally define each task in the following sections.
@@ 2.1.2 Image Scene
$\bullet$ Referring Expression Segmentation (RES). RES, also known as Referring Image Segmentation (RIS), aims to segment the target object in a given image ${I}\in\mathbb{R}^{H\times W\times 3}$ according to a natural language referring expression ${E}$. The output is a binary mask ${M}\in\{0,1\}^{H\times W}$ that precisely delineates the object referred to by ${E}$. Compared to semantic or instance segmentation tasks that rely on predefined object categories, RES introduces new challenges in comprehending the linguistic content of ${E}$ and reasoning about the complex visual relationships among objects in the scene, such as spatial configurations, attributes, and interactions.
@@ 2.1.3 Video Scene
[…]
@@ 4.7 Other Task Settings
$\bullet$ Other. Beyond the core tasks, alternative settings have been explored [246, 247, 248]. Zero-guidance segmentation [246] seeks to segment input images and label all segments using natural language. Grounded Conversation Generation [247] aims to generate natural language responses interleaved with object segmentation masks. Several methods [249, 250] extend referring segmentation capabilities to embodied intelligence settings. Additionally, PTQ4RIS [251] proposes a post-training quantization framework to address challenges for on-device RES inference.
@@ 6.2 Referring Audio-Visual Segmentation
Ref-AVS [93] extends referring segmentation to the audio-visual domain, aiming to segment specific objects in video using synchronized audio and free-form language expressions. To support this task, Wang et al. [93] introduce the first Ref-AVS benchmark with over 3,000 videos annotated with masks and text expressions. The proposed framework employs modality-specific encoders for audio, vision, and text, fused via a query-based decoder. Ref-AVS handles scenarios beyond the scope of traditional AVS or RVOS, such as “person singing bass in a cappella group”, by leveraging multimodal cues for fine-grained disambiguation. It also shows that adding language improves the discriminative power of audio cues [23]. Omni-R1 [325] further enhances performance on Ref-AVS using reinforcement learning to select keyframes and rewrite tasks, enabling efficient Ref-AVS with just one training epoch. STBridge [292] bridges the modality gap between speech and text, allowing RVOS models to effectively process noisy spoken input.
@@ 6.3 Omnimodal Referring Audio-Visual Segmentation
7 3D Referring Expression Segmentation
Compared to its 2D counterpart, 3D Referring Expression Segmentation (3D-RES) poses new challenges due to the intrinsic complexity of point cloud data. While 2D RES operates on structured, grid-like image pixels, 3D-RES involves identifying semantically relevant points within unordered, irregular, and sparse 3D point clouds. This requires not only effective multimodal alignment between language and visual features, but also a strong understanding of underlying geometric structures. The demand for precise point-wise segmentation and fine-grained semantic reasoning makes 3D-RES a particularly challenging yet critical direction for advancing multimodal understanding in 3D space.
7.1 Fully Supervised Learning
@@
7.2 Weakly Supervised/Unsupervised Methods
@@ A.1 RES Performance Benchmarking
[…]
@@ A.2 RVOS Performance Benchmarking
[…]
@@
$\bullet$ $&\mathcal{J}\&\mathcal{F}$ is computed as the average of region similarity and contour accuracy: $\mathcal{J}\&\mathcal{F}=(\mathcal{J}+\mathcal{F})/2$, providing a comprehensive evaluation of both region and boundary accuracy.
RVOS Benchmark Results. As shown in TABLE III, we present the performance comparison of recent RVOS methods on three widely-used benchmarks: MeViS [1], Ref-YouTube-VOS [38], and Ref-DAVIS17 [34]. The field has witnessed substantial progress since 2020, with recent methods like GLUS [290] and VRS-HQ [113] achieving remarkable improvements on all benchmarks. On the MeViS dataset, which focuses on temporal motion understanding, GLUS achieves the best performance with 51.30% $\mathcal{J}$$&\mathcal{F}$, followed closely by VRS-HQ with 50.90%. For Ref-YouTube-VOS, VRS-HQ leads with 71.00% $\mathcal{J}$$&\mathcal{F}$, significantly outperforming earlier methods like URVOS [38] (47.23%). On Ref-DAVIS17, VRS-HQ also demonstrates superior performance with 74.40% $\mathcal{J}$$&\mathcal{F}$, showing substantial improvement over previous state-of-the-art methods.
@@ A.3 R-AVS Performance Benchmarking
[…]
@@
[…]
@@
TABLE VI: Performance Comparison on OmniAVS Benchmark. The results are evaluated on OmniAVS [8] dataset with $&\mathcal{J}\&\mathcal{F}$ as the default metric. All is the average result across 8 splits. MET.: METEOR.
@@
AVS Benchmark Results. As shown in TABLE IV, we present the performance comparison of recent AVS methods on AVSBench [41] (S4 and MS3 subsets) and AVSBench-Semantic [42] (AVSS) using $\mathcal{J}$ and $\mathcal{F}$ metrics. The field has seen significant progress since 2022, with recent methods like DDESeg [313] and RAVS [312] achieving remarkable improvements on all benchmarks. DDESeg achieves the best performance on S4 with 92.40% $\mathcal{J}$ and $\mathcal{F}$, as well as on MS3 with 72.30% $\mathcal{J}$ and $\mathcal{F}$. For the more challenging AVSS benchmark, DDESeg also leads with 63.40% $\mathcal{J}$ and $\mathcal{F}$, demonstrating the effectiveness in audio-visual semantic segmentation task.
Ref-AVS Benchmark Results. As shown in TABLE V, we present the performance comparison of recent Ref-AVS methods on the Ref-AVS [93] dataset using $\mathcal{J}$ and $\mathcal{F}$ metrics across different scenarios. The field has seen remarkable progress since 2022, with recent methods like OmniAVS [8] achieving the best overall performance with 51.7% $\mathcal{J}$ and $\mathcal{F}$ on the Seen set, and 58.3% $\mathcal{J}$ and $\mathcal{F}$ on the Unseen set. TSAM [373] and SAM2-LOVE [374] also demonstrate competitive results, with TSAM achieving 43.4% $\mathcal{J}$ and $\mathcal{F}$ on Seen, and SAM2-LOVE reaching 66.5% $\mathcal{J}$ and $\mathcal{F}$ on Unseen scenarios. These results highlight the significant advancement in audio-visual referring segmentation capabilities, substantially outperforming earlier methods like AVSBench [41] and ReferFormer [271].
OmniAVS Benchmark Results. As shown in TABLE VI, we present the performance comparison of recent methods on the OmniAVS [8] dataset using $\mathcal{J}$$&\mathcal{F}$ metric across eight different splits and METEOR for caption evaluation. OmniAVS [8] achieves the best overall performance with 40.5% $\mathcal{J}$$&\mathcal{F}$ across all splits and 17.0 METEOR score, demonstrating significant improvements over previous methods. MUTR [278] and LISA-13B [19] also show competitive results with 35.2% and 35.4% overall performance respectively. The results across different splits show varying levels of difficulty, with splits VII and VIII generally achieving higher performance (48.2-53.6%) compared to splits III and IV (26.9-36.2%), indicating the heterogeneous nature of the benchmark and the challenges in multi-modal audio-visual understanding.
@@
Evaluation Metrics. IoU and Acc@X (i.e., Pr@X) are adopted for evaluating 3D-RES performance.
TABLE VII: Performance Comparison on 3D-RES Benchmarks. The results are evaluated on ScanRefer [43] dataset with Acc@K metric. denotes weakly-supervised methods.
Method Venue Unique ($\sim$19%) Multiple ($\sim$81%) Overall
@@
TGNN [58] [AAAI’21] - - - - 37.50 31.40
X-RefSeg3D [326] [AAAI’24] - - - - 40.33 33.77
3D-STMN [327] [AAAI’24] 89.30 84.00 46.20 29.20 54.60 39.80
RefMask3D [74] [MM’24] 89.55 84.69 48.09 40.77 55.87 49.24
MDIN [9] [MM’24] 91.00 87.20 50.10 44.90 58.00 53.10
MCLN [336] [ECCV’24] 89.57 78.22 53.28 45.88 58.70 50.70
RG-SAN [330] [NeurIPS’24] 89.20 84.30 55.00 35.40 61.70 44.90
UniSeg3D [338] [NeurIPS’24] - - - - 41.50 28.00
LESS [328] [NeurIPS’24] - - - - 53.23 29.88
Reason3D [332] [3DV’25] 88.40 84.20 50.50 31.70 57.90 41.90
IPDN [331] [AAAI’25] 91.50 88.00 53.10 47.00 60.60 54.90

### v2

@@ 1 Introduction
Multimodal referring segmentation [1, 2, 3, 4, 5, 6, 7, 8, 9] aims to segment the target object in a visual scene of image [3, 4, 7], video [1, 2, 10], or 3D [11, 5, 9] according to a referring expression, such as free-form text or audio. For example, as shown in Fig. 1(b), given the text referring expression “The bird flying away”, the model is expected to segment and track the described target object in the video. This task presents a fundamental and challenging problem in multimodal understanding and supports a wide range of practical applications, such as image/video editing [12, 13], robotics [14], autonomous driving [15], etc. Because of its significant potential in practical applications, multimodal referring segmentation has received growing attention in recent years, as shown in Fig. 3.
@@ 2.1.1 A Unified Formulation
where $f$ is typically instantiated as a neural network. The input space $\mathcal{X}$ consists of two components: the visual input $\mathcal{V}$ (e.g., image, silent video, auditory video, or 3D data), and the referring signal E{E} (e.g., text, audio, etc.) that specifies the target object(s) of interest. The output space $\mathcal{Y}$ consists of segmentation mask(s) of the referred entities within $\mathcal{V}$. Based on this unified formulation, we construct a comprehensive taxonomy of referring segmentation tasks and formally define each task in the following sections.
@@ 2.1.2 Image Scene
$\bullet$ Referring Expression Segmentation (RES). RES, also known as Referring Image Segmentation (RIS), aims to segment the target object in a given image ${I}\in\mathbb{R}^{H\times W\times 3}$ according to a natural language referring expression E{E}. The output is a binary mask ${M}\in\{0,1\}^{H\times W}$ that precisely delineates the object referred to by E{E}. Compared to semantic or instance segmentation tasks that rely on predefined object categories, RES introduces new challenges in comprehending the linguistic content of E{E} and reasoning about the complex visual relationships among objects in the scene, such as spatial configurations, attributes, and interactions.
@@
MeViSv2 [2] 2,006 33,072 8,171 Based on MeViS, further adding no-target and motion reasoning expressions, adding audio expressions.
@@
Ref-LERF [5] 4 295 59 Focusing on spatial relationships for referring 3D gaussian splatting segmentation.
@@ 2.1.3 Video Scene
$\bullet$ Audio-Visual Segmentation (AVS). AVS aims to segment the sound-emitting objects throughout an auditory video. Given an auditory video $\{\textbf{V},\textbf{A}\}$, where $\textbf{V}=\{V_{t}\}_{t=1}^{T}$ denotes $T$ visual frames and $\textbf{A}=\{A_{t}\}_{t=1}^{T}$ is the corresponding audio stream, the goal of AVS is to predict binary masks $\textbf{M}=\{M_{t}\}_{t=1}^{T}$, with each $M_{t}\in{0,1}^{H\times W}$ highlighting the regions in $V_{t}$ associated with the sound source in $A_{t}$. AVS is a special case of referring video segmentation, where the query is implicitly defined as “segment the sound-emitting objects in the video.”
$\bullet$ Referring Audio-Visual Segmentation (Ref-AVS). Ref-AVS aims to segment target objects in an auditory video $\{\textbf{V},\textbf{A}\}$ according to a text referring expression $E$. The desired output is a sequence of binary masks $\textbf{M}=\{M_{t}\}_{t=1}^{T}$, where each $M_{t}\in\{0,1\}^{H\times W}$ is the pixel-level segmentation of the referred object in frame $V_{t}$. Ref-AVS enables handling scenarios that are difficult to address in AVS and RVOS, e.g., “segment the person singing bass in the a cappella group”. This poses the importance of leveraging multi-modal cues to guide visual segmentation.
$\bullet$ Omnimodal Referring Audio-Visual Segmentation (OmniAVS). OmniAVS aims to segment specific target objects in an auditory video $\{\textbf{V},\textbf{A}\}$ according to a multimodal expression $E$ that flexibly combines text, speech, sound, and visual cues. The desired output is a sequence of binary masks $\textbf{M}=\{M_{t}\}_{t=1}^{T}$, where each $M_{t}\in\{0,1\}^{H\times W}$ denotes the pixel-level segmentation of the referred object in frame $V_{t}$. The ability to handle diverse multimodal expressions makes OmniAVS both practical for real-world applications and well-suited for advancing omnimodal models with fine-grained perceptual capabilities.
@@ 2.2.2 Video Scene
$\bullet$ MeViSv2 [2] extends MeViS [1] with 4,502 additional motion reasoning and no-target expressions, and further provides audio expressions to support audio-based RVOS.
@@ 2.2.3 Auditory Video Scenes
$\bullet$ Ref-LERF [5] consists of 772 training images and 22 testing images sourced from LERF [59], and is newly annotated with 295 language expressions for 59 objects. The expressions focus on spatial relationships, providing rich contextual cues for referring segmentation in 3D Gaussian Splatting.
@@ 4.7 Other Task Settings
$\bullet$ Other. Beyond the core tasks, alternative settings have been explored [249, 250, 251, 252]. Zero-guidance segmentation [249] seeks to segment input images and label all segments using natural language. Grounded Conversation Generation [250] aims to generate natural language responses interleaved with object segmentation masks. Several methods [253, 254] extend referring segmentation capabilities to embodied intelligence settings. Additionally, PTQ4RIS [255] proposes a post-training quantization framework to address challenges for on-device RES inference.
@@ 6.2 Referring Audio-Visual Segmentation
Ref-AVS [96] extends referring segmentation to the audio-visual domain, aiming to segment specific objects in video using synchronized audio and free-form language expressions. To support this task, Wang et al. [96] introduce the first Ref-AVS benchmark with over 3,000 videos annotated with masks and text expressions. The proposed framework employs modality-specific encoders for audio, vision, and text, fused via a query-based decoder. Ref-AVS handles scenarios beyond the scope of traditional AVS or RVOS, such as “person singing bass in a cappella group”, by leveraging multimodal cues for fine-grained disambiguation. It also shows that adding language improves the discriminative power of audio cues [25]. Omni-R1 [329] further enhances performance on Ref-AVS using reinforcement learning to select keyframes and rewrite tasks, enabling efficient Ref-AVS with just one training epoch. STBridge [296] bridges the modality gap between speech and text, allowing RVOS models to effectively process noisy spoken input. SAM2-LOVE [330] introduces a multimodal fusion framework that integrates audio, text, and visual information into learnable tokens to prompt SAM2 [128] for Ref-AVS.
@@ 6.3 Omnimodal Referring Audio-Visual Segmentation
7 Referring 3D Segmentation
7.1 3D Referring Expression Segmentation
7.1.1 Fully Supervised Learning
@@
7.1.2 Weakly Supervised/Unsupervised Methods
@@
7.2 Referring 3D Gaussian Splatting Segmentation
To advance research in Referring 3D Gaussian Splatting Segmentation (R3DGS), ReferSplat [5] introduces a framework that links 3D Gaussians with language via spatially aware referring fields, and proposes Ref-LERF, the first dedicated dataset for this task. Each Gaussian is assigned a learnable feature vector that interacts with textual queries to produce segmentation masks through similarity-based rendering.
Despite recent progress, several challenges remain in R3DGS. First, existing method [5] is limited to static scenes and lack temporal modeling. Integrating 4D representations, e.g., 4D Gaussian Splatting [346], could enable reasoning in dynamic environments. Second, current method emphasizes segmentation but overlook fine-grained grounding, such as precise localization and size estimation, which are critical for spatial reasoning. Finally, the limited scale and diversity of existing dataset restrict generalization. Advancing the field requires the development of large-scale, diverse datasets to bridge the gap with more mature 2D counterparts.
@@ A.1 RES Performance Benchmarking
TABLE II: Performance Comparison on RES benchmarks. The results are evaluated on RefCOCO [8], RefCOCO+ [8], and RefCOCOg [374] (UMD partition) using mIoU. $\dagger$: results post-processed with DenseCRF [375]. *: methods utilizing 30% mask annotations and 70% bounding box annotations. ‘ft’ denotes fine-tuning on referring expression segmentation datasets. $\ddagger$} indicates methods using only 5% labeled data. UNINEXT [194] utilizes val/test set masks during training, leading to mask information leakage.
@@
$\mathcal{J}$$\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}$$\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}$$\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$
@@
Sa2VA-26B [122] [ArXiv’25] 46.20 - - 70.10 - - 77.00 - -
RGA3-7B [376] [ICCV’25] 50.10 47.40 52.80 68.50 66.80 70.10 72.80 68.30 77.30
MPG-SAM 2 [279] [ICCV’25] 53.70 50.70 56.70 73.90 71.70 76.10 72.40 68.80 78.00
A.2 RVOS Performance Benchmarking
@@
$\bullet$ $\mathcal{J}\&\mathcal{F}$ is computed as the average of region similarity and contour accuracy: $\mathcal{J}\&\mathcal{F}=(\mathcal{J}+\mathcal{F})/2$, providing a comprehensive evaluation of both region and boundary accuracy.
RVOS Benchmark Results. As shown in TABLE III, we present the performance comparison of recent RVOS methods on three widely-used benchmarks: MeViS [1], Ref-YouTube-VOS [40], and Ref-DAVIS17 [36]. The field has witnessed substantial progress since 2020, with recent methods like GLUS [294] and VRS-HQ [116] achieving remarkable improvements on all benchmarks. On the MeViS dataset, which focuses on temporal motion understanding, GLUS achieves the best performance with 51.30% $\mathcal{J}$$\mathcal{F}$, followed closely by VRS-HQ with 50.90%. For Ref-YouTube-VOS, VRS-HQ leads with 71.00% $\mathcal{J}$$\mathcal{F}$, significantly outperforming earlier methods like URVOS [40] (47.23%). On Ref-DAVIS17, VRS-HQ also demonstrates superior performance with 74.40% $\mathcal{J}$$\mathcal{F}$, showing substantial improvement over previous state-of-the-art methods.
@@ A.3 R-AVS Performance Benchmarking
$\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$
@@
$\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ $\mathcal{S}$ ($\downarrow$)
@@
TABLE VI: Performance Comparison on OmniAVS Benchmark. The results are evaluated on OmniAVS [10] dataset with $\mathcal{J}\&\mathcal{F}$. All is the average result across 8 splits. MET.: METEOR.
@@
AVS Benchmark Results. As shown in TABLE IV, we present the performance comparison of recent AVS methods on AVSBench [43] (S4 and MS3 subsets) and AVSBench-Semantic [44] (AVSS) using $\mathcal{J}$ and $\mathcal{F}$ metrics. The field has seen significant progress since 2022, with recent methods like DDESeg [317] and RAVS [316] achieving remarkable improvements on all benchmarks. DDESeg achieves the best performance on S4 with 92.40% $\mathcal{J}$ and 95.90% $\mathcal{F}$, as well as on MS3 with 72.30% $\mathcal{J}$ and 83.40% $\mathcal{F}$. For the more challenging AVSS benchmark, DDESeg also leads with 63.40% $\mathcal{J}$ and 72.30% $\mathcal{F}$, demonstrating the effectiveness in audio-visual semantic segmentation task.
Ref-AVS Benchmark Results. As shown in TABLE V, we present the performance comparison of recent Ref-AVS methods on the Ref-AVS [96] dataset using $\mathcal{J}$ and $\mathcal{F}$ metrics across different scenarios. The field has seen remarkable progress since 2022, with recent methods like OmniAVS [10] achieving the best overall performance with 51.7% $\mathcal{J}$ and 58.7% $\mathcal{F}$ on the Seen set, and 58.3% $\mathcal{J}$ and 65.1% $\mathcal{F}$ on the Unseen set. TSAM [380] and SAM2-LOVE [330] also demonstrate competitive results, with TSAM achieving 43.4% $\mathcal{J}$ and 56.8% $\mathcal{F}$ on Seen, and SAM2-LOVE reaching 66.5% $\mathcal{J}$ and 72.3% $\mathcal{F}$ on Unseen scenarios. These results highlight the significant advancement in audio-visual referring segmentation capabilities, substantially outperforming earlier methods like AVSBench [43] and ReferFormer [275].
OmniAVS Benchmark Results. As shown in TABLE VI, we present the performance comparison of recent methods on the OmniAVS [10] dataset using $\mathcal{J}$$\mathcal{F}$ metric across eight different splits and METEOR for caption evaluation. OISA-1B [10] achieves the best overall performance with 41.1% $\mathcal{J}$$\mathcal{F}$ across all splits and 21.7 METEOR score, demonstrating significant improvements over previous methods. MUTR [282] and LISA-13B [21] also show competitive results with 32.3% and 36.1% overall performance respectively. The results across different splits show varying levels of difficulty, with splits VII and VIII generally achieving higher performance (e.g., OISA-1B’s 52.6% on VII) compared to splits III and IV (e.g., OISA-1B’s 34.9% on III), indicating the heterogeneous nature of the benchmark and the challenges in multi-modal audio-visual understanding.
@@
Evaluation Metrics. mIoU and Acc@X (i.e., Pr@X) are adopted for evaluating 3D-RES performance.
TABLE VII: Performance Comparison on 3D-RES Benchmarks. The results are evaluated on ScanRefer [45] dataset with Acc@K and mIoU metric. denotes weakly-supervised methods.
Method Venue Unique ($\sim$19%) Multiple ($\sim$81%) Overall mIoU
@@
TGNN [61] [AAAI’21] - - - - 37.50 31.40 28.80
X-RefSeg3D [331] [AAAI’24] - - - - 40.33 33.77 29.94
3D-STMN [332] [AAAI’24] 89.30 84.00 46.20 29.20 54.60 39.80 39.50
RefMask3D [77] [MM’24] 89.55 84.69 48.09 40.77 55.87 49.24 44.86
MDIN [11] [MM’24] 91.00 87.20 50.10 44.90 58.00 53.10 48.30
SegPoint [9] [ECCV’24] - - - - - - 41.70
MCLN [341] [ECCV’24] 89.57 78.22 53.28 45.88 58.70 50.70 44.72
RG-SAN [335] [NeurIPS’24] 89.20 84.30 55.00 35.40 61.70 44.90 44.60
UniSeg3D [343] [NeurIPS’24] - - - - 41.50 28.00 29.60
LESS [333] [NeurIPS’24] - - - - 53.23 29.88 33.74
Reason3D [337] [3DV’25] 88.40 84.20 50.50 31.70 57.90 41.90 42.00
3D-LLaVA [339] [CVPR’25] - - - - - - 43.30
IPDN [336] [AAAI’25] 91.50 88.00 53.10 47.00 60.60 54.90 50.20